OpenStudy (scarlettfarra2000):

Which value of m makes the statement true? 83-2m≤8+3m A) 4 B) 7 C) 12 D) 15

zepdrix (zepdrix):

\[\large\rm 83-2(\color{orangered}{m})\le 8+3(\color{orangered}{m})\]Hmm let's try plugging in \(\large\rm \color{orangered}{m=4}\) and see what happens.

zepdrix (zepdrix):

\[\large\rm 83-2(\color{orangered}{4})\le 8+3(\color{orangered}{4})\]

zepdrix (zepdrix):

Multiplying some stuff, and then doing addition/subtraction where needed, gives us,\[\large\rm 75\le 20\]75 is NOT less than or equal to 20. Ooo looks like m=4 is not the one we're looking for.

zepdrix (zepdrix):

Can you try one of the other options using the same steps? :)

OpenStudy (scarlettfarra2000):

\[83-2(12)\le8+3(12)\]

OpenStudy (scarlettfarra2000):

\[83-24\le8+36\]